OpenStudy (anonymous):

Check my answer! Solve for log3x+4 4096=4 My answer is x = 0. Is this right?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I believe that the equation is: \[\log_{3x+4}4096 =4\]

OpenStudy (anonymous):

Thats correct,

OpenStudy (anonymous):

The question asked to solve for x, the choices were x = -1 x = 0 x = 4/3 x = 3. I chose 0. Is this correct? @DominicNg

OpenStudy (anonymous):

Hi, incorrect. Should be 4/3.

OpenStudy (anonymous):

That makes more sense, thank you!